# Bimanual Manipulation of Steady Hand Eye Robots with Adaptive Sclera Force Control: Cooperative vs. Teleoperation Strategies

### v1

@@ IV-A Sclera Force Control Performance
Therefore, based on the observed results, the adaptive bimanual teleoperation control mode demonstrated the desired performance during the vessel-following experiment and could be a promising technology with new capabilities to improve surgeons’ comfort, dexterity, and patients’ safety. The effectiveness of the proposed BMAT control mode should be further examined using other types of biological models, such as chicken embryo or pig eye, in addition to this phantom eye model. Moreover, adding haptic force feedback to the teleoperation mode provides the surgeons with a better understanding of the sclera force level and may improve the safety of surgery.
V Conclusion
@@
This work performed a vessel-following experiment on a bimanual adaptive teleoperation framework with an adaptive sclera force control algorithm on SHER 2.0 and SHER 2.1 using two force-sensing tools. This was a first in robot-assisted retinal surgery applications. The tools were equipped with FBG sensors capable of measuring the force at the tool sclerotomy and the tooltip. The performance of this bimanual teleoperation mode was compared with a bimanual cooperative mode utilizing a single expert user by doing a vessel-following experiment inside an eye phantom for two sitting and standing postures. Both surgical tools, including the secondary tool used for rotating the eyeball, are force-sensing tools that provide sclera force measurements for the adaptive force control algorithm to adjust scleral forces automatically to maintain them below a safe threshold. Using this force control technique, it is possible to perform bimanual robotic surgery on the eye without registering the two robots.
Also, a comparison between the bimanual teleoperation and the bimanual cooperative modes in the sitting posture demonstrates that the bimanual teleoperation mode shows comparable performance to the bimanual cooperative mode and can potentially improve the safety and precision of surgery by providing surgeons with additional capabilities, including motion scaling and repositioning.
In the future, we intend to add haptic feedback to reflect the sclera force on the user’s hand and conduct a multi-user study to further compare the BMAT control mode equipped with haptic feedback and the BMAC control mode.

### v2

@@ V Discussion and Future Work
Based on the observed results, the proposed BMAT control mode demonstrated the desired performance during the vessel-following experiment. It could be a promising technology with new capabilities to improve surgeons’ comfort, dexterity, and patients’ safety.
One limitation of this work is that the experiments are conducted by a single user. This is a pilot study to analyze the performance of the different control modes. Moreover, since the force-sensing tools are fragile and expensive to build, it is important to work with expert users with enough training in bimanual teleoperation control of the robots. For future work, the effectiveness of the proposed BMAT control mode should be further examined using other types of biological models, such as chicken embryo or pig eye. Moreover, integrating this BMAT mode with haptic force feedback [28, 29, 30, 31], augmented [32] or mixed reality [33] technology, and learning from demonstration (LfD) [34, 35] may improve non-experts’ ability to efficiently manipulate and control the robots [36] and improve the safety of surgery.
VI Conclusion
@@
This work performed a vessel-following experiment on a bimanual adaptive teleoperation framework with an adaptive sclera force control algorithm on SHER 2.0 and SHER 2.1 using two FBG-based force-sensing tools. The tools were equipped with FBG sensors capable of measuring the force at the tool sclerotomy and the tooltip. The performance of this bimanual teleoperation mode was compared with a bimanual cooperative mode utilizing a single expert user by doing a vessel-following experiment inside an eye phantom for two sitting and standing postures. Using this force control technique, it is possible to safely perform bimanual robotic manipulation of the eye without registering the two robots. Also, a comparison between the BMAT and the BMAC modes in the sitting posture demonstrates that the BMAT mode shows comparable performance to the BMAC mode and can potentially improve the safety and precision of surgery by providing surgeons with additional capabilities, including motion scaling and repositioning.